### v1

Abstract
Arrival is the problem of deciding which out of two possible destinations will be reached first by a token that moves deterministically along the edges of a directed graph, according to so-called switching rules. It is known to lie in $\mathsf{NP}\cap\mathsf{CoNP}$, but not known to lie in $\mathsf{P}$. The state-of-the-art algorithm due to Gärtner et al. (ICALP ‘21) runs in time $2^{\mathcal{O}(\sqrt{n}\log n)}$ on an $n$-vertex graph.
We prove that Arrival can be solved in time $2^{\mathcal{O}(k\log^{2}n)}$ on $n$-vertex graphs of treewidth $k$. Our algorithm is derived by adapting a simple recursive algorithm for a generalization of Arrival called G-Arrival. This simple recursive algorithm acts as a framework from which we can also rederive the subexponential upper bound of Gärtner et al.
Our second result is a reduction from G-Arrival to the problem of finding an approximate fixed point of an $\ell_{1}$-contracting function $f:[0,1]^{n}\rightarrow[0,1]^{n}$. Finding such fixed points is a well-studied problem in the case of the $\ell_{2}$-metric and the $\ell_{\infty}$-metric, but little is known about the $\ell_{1}$-case.
Both of our results highlight parallels between Arrival and the Simple Stochastic Games (SSG) problem. Concretely, Chatterjee et al. (SODA ‘23) gave an algorithm for SSG parameterized by treewidth that achieves a similar bound as we do for Arrival, and SSG is known to reduce to $\ell_{\infty}$-contraction.
keywords:
ARRIVAL, G-ARRIVAL, Deterministic Random Walk, Rotor-Routing, $\ell_{1}$-Contraction, Banach Fixed Point.
category:
[…]
1 Introduction
Arrival is a computational problem first introduced by Dohrau et al. [dohrauARRIVALZeroPlayerGraph2017]. It can be described as a deterministic process (or zero-player game) on a directed graph with a designated origin $o$ and two designated destinations $d$ and $\overline{d}$. Every vertex in Arrival has out-degree two, and exactly one outgoing edge at every vertex is marked (we also call it the even edge). We additionally assume that both destinations are reachable from every vertex in the graph. A token is placed on $o$ and moved along the edges of the graph according to the following rule: At every vertex, the token continues along the outgoing edge that was used least so far. In case of a tie, the token uses the even edge. This effectively means that the token will alternate between the two outgoing edges at every vertex, starting with the even edge. The task is to decide which of the two destinations $d$ or $\overline{d}$ will be visited first by the token.
Dohrau et al. [dohrauARRIVALZeroPlayerGraph2017] proved that Arrival is contained in $\mathsf{NP}\cap\mathsf{CoNP}$. Naturally, they then asked whether it is also in $\mathsf{P}$. This open problem has received some attention in recent years and the best algorithm to date, due to Gärtner et al. [gartnerSubexponentialAlgorithmARRIVAL2021], runs in time $2^{\mathcal{O}(\sqrt{n}\log n)}$ on a graph with $n$ vertices.
We present two new results for Arrival: Our first result is an algorithm for Arrival that runs in time $2^{\mathcal{O}\left(k\log^{2}n\right)}$ on graphs with $n$ vertices and treewidth $k$. Note that this bound is quasi-polynomial for graphs with bounded treewidth. Our algorithm is obtained by adapting a simple recursive algorithm for G-Arrival, a generalization of Arrival that allows arbitrarily many origins, destinations, and tokens: The recursive algorithm solves an instance with $\ell$ destinations and origins by making recursive calls on instances with $\ell+1$ destinations and origins. In other words, in each recursive call, a new vertex is made into a destination and origin vertex. By choosing this pivot vertex carefully, we can exploit the underlying graph structure.
It turns out that this simple recursive algorithm for G-Arrival can also be seen as a framework for other algorithms for G-Arrival. Concretely, we explain how the state-of-the-art upper bound $2^{\mathcal{O}(\sqrt{n}\log n)}$ due to Gärtner et al. [gartnerSubexponentialAlgorithmARRIVAL2021] can be derived in this framework as well.
Our second result is a reduction from G-Arrival to the problem of finding an approximate fixed point of a $\ell_{1}$-contracting function $f:[0,1]^{n}\rightarrow[0,1]^{n}$. Concretely, we say that a function $f:[0,1]^{n}\rightarrow[0,1]^{n}$ is contracting with parameter $\lambda\in[0,1)$ if we have $\lVert f(x)-f(y)\rVert_{1}\leq\lambda\lVert x-y\rVert_{1}$ for all $x,y\in[0,1]^{n}$. Such a function is guaranteed to have a unique fixed point by Banach’s fixed point theorem [banach1922operations]. An $\varepsilon$-approximate fixed point $x\in[0,1]^{n}$ has to satisfy $\lVert f(x)-x\rVert_{1}\leq\varepsilon$. Given our reduction, any algorithm that can find an $\varepsilon$-approximate fixed point of $f$ in time $\operatorname{poly}\left(\log\frac{1}{\varepsilon},\log\frac{1}{1-\lambda},n\right)$ would imply a polynomial-time algorithm for G-Arrival.
Finding approximate fixed points is a well-studied problem in the case of the $\ell_{2}$-metric and the $\ell_{\infty}$-metric (see e.g. [sikorskiComputationalComplexityFixed2009, chenComputingFixedPoint2024]). In particular, efficient algorithms in the case of the $\ell_{2}$-metric have been known since 1993 [sikorskiEllipsoidAlgorithmComputation1993], and only recently the first polynomial query upper bound for the $\ell_{\infty}$-metric was found by Chen et al. [chenComputingFixedPoint2024]. They also mention the case of the $\ell_{1}$-metric as an open problem, albeit without mentioning any applications. To the best of our knowledge, our reduction from G-Arrival to $\ell_{1}$-contraction provides the first concrete application for this problem, further motivating the study of its complexity. We think that our reduction could eventually lead to better algorithms for Arrival through the study of $\ell_{1}$-contraction maps.
1.1 Related Work
Since Arrival is contained in $\mathsf{NP}\cap\mathsf{CoNP}$, it also naturally fits into the complexity class $\mathsf{TFNP}$, which contains total problems with efficiently verifiable solutions. In fact, after a series of results for containment in subclasses of $\mathsf{TFNP}$ [karthikc.s.DidTrainReach2017, gartnerARRIVALNextStop2018], we now know that Arrival is contained in $\mathsf{UEOPL}$ [fearnleyUniqueEndPotential2020].
Going into a slightly different direction, Gärtner et al. [gartnerARRIVALNextStop2018] proved that Arrival is also contained in $\mathsf{UP}$ and $\mathsf{CoUP}$, the analogues of $\mathsf{NP}$ and $\mathsf{CoNP}$ with unique solutions. In fact, it would not be hard to rederive this using our reduction to $\ell_{1}$-contraction: The idea is that the fixed point acts as an efficient certificate for both YES- and NO-instances and it must be unique due to the contraction property.
It is also known that Arrival can be solved in polynomial time on some restricted graph classes. For example, a result due to Priezzhev et al. [priezzhevEulerianWalkersModel1996] implies that Arrival can be solved by simulation in polynomial time on Eulerian graphs. Other results include polynomial-time algorithms on tree-like multigraphs [augerPolynomialTimeAlgorithm2022] and path-like multigraphs with many tokens [augerGeneralizedARRIVALProblem2023].
Finally, further variants of Arrival have been studied in the past, including a stochastic variant [websterStochasticArrivalProblem2022], a recursive variant [websterRecursiveArrivalProblem2023], as well as variants with one or two players [fearnleyReachabilitySwitchingGames2021].
Comparison to SSG
As mentioned before, both our results show parallels between Arrival and Simple Stochastic Games (SSG). We will briefly discuss the connections between the two problems. Similarly to Arrival, SSG is contained in $\mathsf{NP}\cap\mathsf{CoNP}$ [condonComplexityStochasticGames1992] and even $\mathsf{UP}\cap\mathsf{CoUP}$ [chatterjeeReductionParityGames2011a], but no polynomial-time algorithm is known. The state-of-the-art algorithm due to Ludwig [ludwigSubexponentialRandomizedAlgorithm1995] runs in randomized subexponential time $2^{\mathcal{O}(\sqrt{n}\log n)}$.
Dohrau et al. [dohrauARRIVALZeroPlayerGraph2017] already wondered about similarities between Arrival and SSG when they first introduced Arrival. Since then, both problems were shown to reduce to the problem of finding a Tarski fixed point [etessamiTarskiTheoremSupermodular2020, gartnerSubexponentialAlgorithmARRIVAL2021], and to be contained in the complexity class $\mathsf{UEOPL}$ [fearnleyUniqueEndPotential2020]. Both our results for Arrival further extend this list of similarities: Our upper bound of $2^{\mathcal{O}(k\log^{2}n)}$ for Arrival on $n$-vertex graphs of treewidth $k$ is comparable to a similar bound for SSG due to Chatterjee et al. [chatterjeeFasterAlgorithmTurnbased2023]. Moreover, SSG reduces to finding an approximate fixed point of an $\ell_{\infty}$-contracting function [condonComplexityStochasticGames1992], which is analogous to our reduction from Arrival to $\ell_{1}$-contraction.
Both Arrival and SSG also admit polynomial-time algorithms on graphs with a bounded feedback vertex set [gartnerSubexponentialAlgorithmARRIVAL2021, augerFindingOptimalStrategies2014]. In fact, in the case of Arrival, we will discuss this in more detail in Section LABEL:ssec:subexponential_algo.
1.2 Outline
As explained above, our results are actually obtained for a generalization of Arrival called G-Arrival. Thus, we will start Section 2 by formally introducing G-Arrival. We also use Section 2 to recall further terminology and notation from the literature that will be useful for our arguments.
Note that instead of formally introducing the notion of treewidth and tree decompositions, we will directly work with so-called balanced separators instead. The reason for this choice of exposition is that our parameterized algorithm actually exploits the existence of small balanced separators (and not tree decompositions themselves), which are guaranteed to exist in graphs of small treewidth (see Section 2.1 for more details).
We describe our parameterized algorithm in Section LABEL:sec:algorithm. We start the exposition with our simple recursive algorithm for G-Arrival (Section LABEL:ssec:recursive_algorithm), which provides a framework from which we will derive the parameterized algorithm in Section LABEL:ssec:exploiting_balanced_separators. In Section LABEL:ssec:subexponential_algo, we additionally explain how the subexponential upper bound due to Gärtner et al. [gartnerSubexponentialAlgorithmARRIVAL2021] and polynomial-time upper bounds on graphs with a bounded feedback vertex set [gartnerSubexponentialAlgorithmARRIVAL2021] can be rederived from our framework.
Finally, Section LABEL:sec:solving_arrival_via_contraction contains our reduction to the problem of finding an approximate fixed point of a $\ell_{1}$-contraction map. Crucially, given an instance of G-Arrival, we define a function $f:\mathbb{R}^{n}_{\geq 0}\rightarrow\mathbb{R}^{n}_{\geq 0}$ that is contracting and thus has a unique fixed point. We then prove that a reasonably good approximation of this fixed point will give away the solution to the G-Arrival-instance. The function can be restricted to a compact subset of $\mathbb{R}^{n}_{\geq 0}$ and scaled to fit into $[0,1]^{n}$, if desired.
2 Preliminaries
We start by recalling G-Arrival, which was first formally defined by Hoang [hoangTwoCombinatorialReconfiguration2022]. Note that our formulation slightly deviates from the one by Hoang, but is easily seen to be equivalent.
A switch graph is a directed graph $G=(V,E,s_{0},s_{1})$ with $s_{0},s_{1}:V\rightarrow V$ and $E=\{(v,s_{0}(v))\mid v\in V\}\cup\{(v,s_{1}(v))\mid v\in V\}$. We consider $E$ to be a multiset, and two edges $(v,s_{0}(v)),(v,s_{1}(v))\in E$ to be distinct objects even if we have $s_{0}(v)=s_{1}(v)$. Given a switch graph, we call $s_{0}(v)$ and $s_{1}(v)$ the even and odd successor of $v\in V$, respectively. Similarly, we refer to edges induced by $s_{0}$ as even edges, and to edges induced by $s_{1}$ as odd edges.
In G-Arrival, many tokens traverse the directed graph simultaneously, starting and ending at special vertices that we call terminals or terminal vertices. Concretely, a G-Arrival-instance consists of a switch graph as well as a non-empty subset $T\subseteq V$ of terminals. For each terminal $v\in T$, we also get a natural number $t^{+}_{v}$ of tokens starting at $v$. We always assume that at least one terminal is reachable from each non-terminal vertex $v\in V\setminus T$ (otherwise, tokens could loop indefinitely without ever reaching a terminal). We will also always denote the total number of tokens by $t^{+}\coloneqq\sum_{v\in T}t^{+}_{v}$ and assume $t^{+}\geq 1$.
Figure 1: An instance of G-Arrival. Even edges are bold while odd edges are dashed. Terminals are marked in red. Three tokens start at the left terminal, and two tokens start at the right terminal.
Now consider the following non-deterministic procedure. Initially, for every terminal $v\in T$, move $\lceil\frac{t^{+}_{v}}{2}\rceil$ tokens from $v$ to $s_{0}(v)$, and $\lfloor\frac{t^{+}_{v}}{2}\rfloor$ tokens from $v$ to $s_{1}(v)$ (we do this for every terminal simultaneously). Then, while there exists a token on a non-terminal $v\in V\setminus T$, non-deterministically choose one such token and move it along the out-edge of $v$ that was used fewer times so far. In case of a tie, the token must use the even out-edge $(v,s_{0}(v))\in E$. In other words, the even and odd out-edges at $v$ will be used in an alternating fashion, starting with the even out-edge. This procedure stops once all tokens have reached a terminal. The goal of G-Arrival is to predict the number of tokens $t^{-}_{v}$ arriving at each terminal $v\in T$.
In order for G-Arrival to be well-defined, we need to make sure that the above procedure terminates and that it always produces the same values $t^{-}_{v}$ for all $v\in T$ (independently of the non-deterministic choices). Indeed, both of these properties hold, and this follows by generalizing the arguments of Dohrau et al. [dohrauARRIVALZeroPlayerGraph2017] to work with multiple tokens, multiple destinations, and multiple origins, as explained by Gärtner et al. [gartnerSubexponentialAlgorithmARRIVAL2021] and Hoang [hoangTwoCombinatorialReconfiguration2022]. To explain this, we first need to recall the concept of switchings flows from the literature.
Given a switch graph $G=(V,E,s_{0},s_{1})$ with terminals $T\subseteq V$ and starting tokens $(t^{+}_{v})_{v\in T}$, a function $x:E\rightarrow\mathbb{R}_{\geq 0}$ satisfying the three constraints
$x(v,s_{0}(v))-x(v,s_{1}(v))$$\in\{0,1\}$ $(\forall v\in V)$
$\underbrace{\sum_{u:(v,u)\in E}x(v,u)}_{=:x^{+}(v)}-\underbrace{\sum_{u:(u,v)\in E}x(u,v)}_{:=x^{-}(v)}$ $=0$ $(\forall v\in V\setminus T)$
$\underbrace{\sum_{u:(v,u)\in E}x(v,u)}_{=:x^{+}(v)}$ $=t^{+}_{v}$ $(\forall v\in T)$
is called a switching flow. We will refer to the first set of constraints above as switching behavior, and to the second set of constraints as flow conservation.
Dohrau et al. [dohrauARRIVALZeroPlayerGraph2017] proved the following theorem in the case of Arrival, but we directly state its generalization for G-Arrival (see also [gartnerSubexponentialAlgorithmARRIVAL2021, hoangTwoCombinatorialReconfiguration2022]).
Theorem 2.1 (Integral Switching Flows are Certificates [dohrauARRIVALZeroPlayerGraph2017]).
Given a switch graph $G=(V,E,s_{0},s_{1})$ with terminals $\varnothing\neq T\subseteq V$ and starting tokens $(t^{+}_{v})_{v\in T}$ with $t^{+}\geq 1$, the number of tokens $t^{-}_{v}$ arriving at terminal $v$ is well-defined and any integral switching flow $x:E\rightarrow\mathbb{N}_{0}$ satisfies $x^{-}(v)=t^{-}_{v}$, for all $v\in T$.
Concretely, Theorem 2.1 states that
Problem 2.2.
G-Arrival is well-defined and that it can be solved by finding any integral switching flow. Observe that by simulating the non-deterministic procedure outlined before and recording the number of times that each edge is traversed by a token, one can obtain a special integral switching flow that we call the run profile. While the run profile is unique (i.e. it does not depend on non-deterministic choices) [gartnerSubexponentialAlgorithmARRIVAL2021], Dohrau et al. [dohrauARRIVALZeroPlayerGraph2017] already observed that in general, integral switching flows are not unique. Concretely, there may be integral switching flows other than the run profile, but Theorem 2.1 tells us that they still predict the values $(t^{-}_{v})_{v\in T}$ correctly.
Figure 2: The number on the edges indicate an integral switching flow. Note that this is not the run profile, but it still certifies (by Theorem 2.1) that four tokens arrive at the right terminal while only one token arrives at the left terminal. To get the run profile, one would have to decrease the flow on the edges of the directed triangle formed by the three non-terminal vertices by one each.
Finally, we will need an upper bound on the total flow in any integral switching flow. Similar upper bounds were used in previous work as well (see e.g. [dohrauARRIVALZeroPlayerGraph2017, gartnerARRIVALNextStop2018, gartnerSubexponentialAlgorithmARRIVAL2021, hoangTwoCombinatorialReconfiguration2022]). We sketch a short proof.
Lemma 2.3 (Upper Bound on Integral Switching Flow [dohrauARRIVALZeroPlayerGraph2017]).
Let $x:E\rightarrow\mathbb{N}_{0}$ be an arbitrary integral switching flow for the
Proof 2.5.
The equation $\sum_{v\in T}t^{+}_{v}=\sum_{v\in T}t^{-}_{v}=\sum_{v\in T}x^{-}(v)$ follows from flow conservation of switching flows and Theorem 2.1. We will now prove the upper bound on the flow values.
Let $e=(u,v)\in E$ be arbitrary. Observe that there must exist a simple path $P=(v_{0},v_{1},v_{2},\dots,v_{k})$ with $v_{0}=v$ of length $0\leq k<|V|$ from $v$ to some $v_{k}\in T$ (recall that we assume that at least one terminal is reachable from every non-terminal in the graph). Observe that by the switching behavior of switching flows, $x(e)\geq 2^{|V|}t^{+}$ would imply $x^{-}(v_{i})\geq 2^{|V|-i}t^{+}$ for all $i\in[k]$. In particular, we would have $t^{-}_{v_{k}}=x^{-}(v_{k})\geq 2t^{+}$, contradicting the equation $t^{+}=\sum_{v\in T}t^{-}_{v}$ above.
2.1 Treewidth and Balanced Separators
Treewidth is a well-established graph parameter that plays an important role in parameterized algorithms, and it is intimately related to the notion of balanced separators (see e.g. [cyganParameterizedAlgorithms2015, Chapter 7]). As is the case with many applications on graphs of bounded treewidth, our algorithm actually exploits the existence of balanced separators and can be formulated without computations of tree decompositions. Hence, we will refrain from formally introducing tree decompositions and instead focus on balanced separators.
Given an undirected graph $G=(V,E)$ and a subset $S\subseteq V$ of its vertices, we use $G-S$ to denote the graph resulting from deleting the vertices in $S$ and their incident edges from $G$. We call $S$ a balanced separator if each connected component in $G-S$ contains at most $\frac{1}{2}|V|$ vertices. Note that this does not necessarily imply that $G-S$ must have more than one connected component and it could even be an empty graph (despite what the term separator may suggest): Concretely, any set $S$ of size at least $\frac{1}{2}|V|$ is a balanced separator, and thus there always exists a balanced separator. Using a brute-force approach, we can find a smallest balanced separator $S$ in $G$ in time $|V|^{\mathcal{O}(|S|)}$.
The following connection between treewidth and balanced separators is crucial for us.
Lemma 2.6 (Balanced Separators and Treewidth [cyganParameterizedAlgorithms2015, Lemma 7.19]).
If $G$ has treewidth at most $k$, then for any $S\subseteq V$, the subgraph $G-S$ has a balanced separator of size at most $k+1$.
In other words, this lemma allows us to use the treewidth of $G$ to infer the existence of small balanced separators in all induced subgraphs of $G$.
All of the previous concepts are defined on undirected graphs. Since we will be working exclusively with directed graphs, we will adopt the following convention: When used on a directed graph, the terms treewidth and balanced separators are to be interpreted with respect to the underlying simple undirected graph.
2.2 Non-Expansive, Contracting, and Monotone Functions
We will be interested in the Manhattan distance, which is induced by the $\ell_{1}$-norm. Concretely, we use $\lVert x\rVert:=\sum_{i=1}^{n}|x_{i}|$ to denote the $\ell_{1}$-norm of a vector $x\in\mathbb{R}^{n}$. The Manhattan distance of $x,y\in\mathbb{R}^{n}$ is then given by $\lVert x-y\rVert$.
We are mainly concerned with the non-negative orthant $\mathbb{R}^{n}_{\geq 0}\subseteq\mathbb{R}^{n}$. A function $f:\mathbb{R}^{n}_{\geq 0}\rightarrow\mathbb{R}^{n}_{\geq 0}$ is called a $\lambda$-contraction (or is $\lambda$-contracting) for some $\lambda\in[0,1)$ if and only if $\lVert f(x)-f(y)\rVert\leq\lambda\lVert x-y\rVert$ for all $x,y\in\mathbb{R}^{n}_{\geq 0}$. Banach’s fixed point theorem [banach1922operations] implies that such a contracting function admits a unique fixed point. If $f$ only satisfies the weaker property $\lVert f(x)-f(y)\rVert\leq\lVert x-y\rVert$ for all $x,y\in\mathbb{R}^{n}_{\geq 0}$, we call it non-expansive instead.
In Section LABEL:sec:solving_arrival_via_contraction, we reduce